OpenStudy (anonymous):

I'm having some trouble on this question, can someone please help, and show the steps? Find 2 numbers whose sum is 17 and whose product is 72.

hartnn (hartnn):

can u factor 72 ?

OpenStudy (anonymous):

What do you mean factor 72? I got up to this step. x+y=17 x*y=72 x=17-y y=17-x (17*y)xy=72 17y-y^2=72 <-------- I'M STUCK ON THIS BIT!!

hartnn (hartnn):

alternate way : use \(\large (x-y)^2 = (x+y)^2-4xy\) to get x-y, then you have x+y and x-y also, which is easily simultaneously solvable.

OpenStudy (anonymous):

x+y=17 x+(17-x) x(17-x)=72 -x^(2)+17x-72=0 x^(2)-17x+72=0 (x-8)(x-9)=0 x=8,9

OpenStudy (anonymous):

Can you please help from the start?

hartnn (hartnn):

sure, you have x+y and xy, right ? you can find x-y from the identity i have given, what u get as x-y from that ?

OpenStudy (anonymous):

let one number be x and the other be y ex. 4+2=6 === 4+(6-4)=6

hartnn (hartnn):

\((x-y)^2 = 17^2-4(72)=...?\)

OpenStudy (anonymous):

then take it from there.:)

OpenStudy (anonymous):

@hartnn you Blow my Mind

hartnn (hartnn):

how exactly ?

hartnn (hartnn):

ohh..literally...

OpenStudy (anonymous):

you r the BEST bro